# Generative Emergent Communication: Large Language Model is a Collective World Model

### v1

@@ Abstract
This study proposes a unifying theoretical framework called generative emergent communication (generative EmCom) that bridges emergent communication, world models, and large language models (LLMs) through the lens of collective predictive coding (CPC). The proposed framework formalizes the emergence of language and symbol systems through decentralized Bayesian inference across multiple agents, extending beyond conventional discriminative model-based approaches to emergent communication. This study makes the following two key contributions: First, we propose generative EmCom as a novel framework for understanding emergent communication, demonstrating how communication emergence in multi-agent reinforcement learning (MARL) can be derived from control as inference while clarifying its relationship to conventional discriminative approaches. Second, we propose a mathematical formulation showing the interpretation of LLMs as collective world models that integrate multiple agents’ experiences through CPC. The framework provides a unified theoretical foundation for understanding how shared symbol systems emerge through collective predictive coding processes, bridging individual cognitive development and societal language evolution. Through mathematical formulations and discussion on prior works, we demonstrate how this framework explains fundamental aspects of language emergence and offers practical insights for understanding LLMs and developing sophisticated AI systems for improving human-AI interaction and multi-agent systems.
Index Terms:
@@
emergent communication, large language model, world model, generative model, symbol emergence.
I Introduction
@@
Langauge evolves and changes over time as a result of decentralized human communications [1, 2, 3, 4]. Sentences are generated to describe a wide range of phenomena, including external events, emotions, and intentions. In particular, the system of language is not static but dynamic [5, 6, 7, 8]. As Peirce, the founder of semiotics, suggested, symbols, including language, can be characterized by a triadic relationship of sign, object, and interpretant [9, 10, 11]. Here, sign corresponds to words, sentences and other signals. In particular, the correspondence between sign and object, which is signified by a sign (i.e., signifier), is determined by an interpretant. In other words, the meaning of a sign, that is, language, depends on culture and context, and so on.
The series of studies on emergent communication (EmCom), which is also referred to as emergent language (EmLang), and symbol emergence has attempted to explain the emergence of language and sharing of meanings of language among agents [2, 12, 13, 14, 15, 16]. However, computational models and general theories that provide a comprehensive and integrative understanding of symbol and language emergence [17, 8] and address the interdependency between the two aspects are lacking: first, world modeling by agents, which depends on their embodiment and environmental adaptation [18, 19]; and second, language emergence, which involves the evolution of language structure reflecting structural knowledge of the world that becomes embedded in language through distributional semantics[20, 21]. Although we are entering the era of large language models (LLMs) and generative artificial intelligence (AI), a need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [22, 8, 23, 16].
Recently, numerous studies have discussed how LLMs possess a model of the world. LLMs learn the distribution of word or linguistic token sequences and become intelligent to the extent that they can solve a wide range of tasks, including question answering, machine translation, and conversations [24, 25]. These capabilities are widely believed to be based on distributional semantics [20, 22]. However, the emergence of a latent structure that enables distributional semantics in our human language is not yet uncovered, though the mechanism through which neural networks find the hidden structure of language is gradually being revealed. Even before the invention of LLMs, word2vec, skip-gram, or continuous bag of words language models could internally form relative knowledge of concepts, for instance, "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [26, 27]. Two language models independently modeling different languages can perform unsupervised machine translation through structural alignment to a certain extent [28]. Recently, it has become clearer that LLMs have much knowledge of the world, including color similarity, physical properties of materials, and spatial knowledge [29, 30, 31, 32, 33]. This implies that the capability of LLMs is based on the mysterious nature of language. We believe that uncovering this nature of human language is the crucial mission of studies on symbol/language emergence and EmCom/EmLang.
In related fields, for instance, evolutionary robotics, cognitive and developmental robotics, artificial life, computational linguistics, and machine learning (ML), a wide range of studies on language evolution, EmCom, and symbol emergence have been conducted [34, 14, 2, 23, 17]. Most of these studies relied on specific types of language games, for instance, referential, signaling, and naming games. Since the late 2010s, empowered by the representation learning and language modeling capability of deep neural networks (DNNs), studies on EmCom have been boosted again [35, 36, 37]. However, most studies only treated the formation of communication protocols in specific game settings.
These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the free-energy principle (FEP), predictive coding (PC), and world modeling. Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 18, 19]. This aligns with broader theoretical frameworks such as PC and the FEP. PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [39], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [40], which is associated with the idea of the Bayesian brain proposed by [41]. Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [42, 43, 44], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
To address these challenges, Taniguchi proposed the collective predictive coding (CPC) hypothesis [16]. The idea is based on a generative view of cognitive systems. The CPC assumes that not only individual agents but also groups of agents committing to symbolic communications can be modeled as generative models. Regarding individual agents, the FEP and active inference provide a generative view of a cognitive system. In other words, the CPC hypothesis extends the idea of the FEP and PC to a societal level [45]. The CPC hypothesis argues that symbol/language emergence can be modeled as decentralized Bayesian inference of a shared latent representation in a hierarchical manner.
This study aims to provide a theoretical view that bridges the idea of world models, that is, internal models for the sensorimotor observations and dynamics, and the emergence of language. To this end, we extend the idea of the CPC hypothesis and formalize the framework of generative emergent communication (generative EmCom) by introducing a generative view to the existing framework of EmCom. This allows us to develop a more integrative general theory of multi-agent systems performing EmCom and symbol emergence systems.
The main contributions of this paper are as follows:
@@
    Propose a new framework of EmCom known as generative EmCom, demonstrating that the emergence of communication in multi-agent reinforcement learning (MARL) can be derived from the perspective of control as inference (CaI) within the framework of generative EmCom, and clarify the relationship between conventional discriminative EmCom and the new generative EmCom.
@@
    Present the mathematical framework showing how LLMs become a collective world model in a specific sense through collective PC.
The remainder of this paper is organized as follows: Section 2 introduces the framework of generative EmCom, formalizing the collective PC hypothesis and its mathematical foundations. Section 3 demonstrates the formulation of language games as decentralized Bayesian inference, with a focus on the Metropolis-Hastings naming game (MHNG). Section 4 explores the application of generative EmCom to MARL, showing how it facilitates cooperative behavior. Section 5 discusses the relationship between the proposed generative framework and conventional approaches to EmCom. Section 6 examines LLMs through the lens of collective world models, providing new insights into their capabilities and limitations. Finally, Section 7 concludes the paper and discusses future directions for research in this area.
II Generative EmCom
@@
Figure 1: (Left) PGM representation of CPC in symbol emergence, that is, emergent communication or language. The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system. The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system. (Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
II-A Collective predictive coding
The CPC hypothesis was proposed to explain the emergence of symbol systems, particularly language [16, 46]. The CPC extends the idea of PC and FEP from the individual cognition level to a societal level. Taniguchi et al. were the first to formulate the CPC from the perspective of FEP and active inference [45]. This suggests that symbol emergence follows a process of minimizing free energy across the entire multi-agent system.
@@ II-A Collective predictive coding
A question is raised. How can we update the external representations, e.g., language, while our brains are disconnected physically? The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., [47, 48]). In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations. A representative example is MHNG explained in Section II-C. The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
@@
Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the collective free energy of a group of agents. This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
II-B Generative EmCom
Numerous studies on EmCom/EmLang have been conducted. Major approaches are language games, MARL, and iterated learning models (ILMs) [12, 6, 49, 50, 51]. Recently, EmCom studies based on language games, particularly referential games, have become a dominant approach in the research community of EmCom. These studies utilize the representation learning capability of DNNs, give agents language encoders and decoders, and make them learn to generate and interpret a complex sequence of tokens to identify its meaning. In a typical setting of referential games, a speaker agent encodes a given object to a sequence of tokens using a neural network, for instance, long short-term memory, and the listener agent decodes it and identifies the object among some other items. This approach considers EmCom as an optimization of communication protocol. Generally, this is in line with Shannon’s communication model [52]. This approach focuses less on the synergy between the perceptual and communicative aspects of language. This approach can be regarded as a discriminative model-based approach to EmCom/EmLang.
Unlike traditional EmCom models, generative EmCom models consider language emergence as a generative model. In the simplest case, a probabilistic graphical model (PGM) of generative EmCom based on CPC hypothesis can be described mathematically as follows:
Generative model:
$p\left(\{x^{k}\}_{k},\{z^{k}\}_{k},m\right)=p(m)\prod_{k}p\left(x^{k}\mid z^{k}\right)p\left(z^{k}\mid m\right)$ (1)
Inference model:
$q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)=q\left(m\mid\{z^{k}\}_{k}\right)\prod_{k}q\left(z^{k}\mid x^{k}\right)$ (2)
where $p(\cdot)$ and $q(\cdot)$ are assumed to be original probability density functions and approximate probability functions, respectively, following the conventional notation of variational inference. Variables and functions used in this paper are listed in Table I. In this framework, $q\left(z^{k}\mid x^{k}\right)$ is interpreted as the process of representation learning performed by agent $k$, whereas $q\left(m\mid\{z^{k}\}_{k}\right)$ is hypothesized to be derived through collaborative language games among the agents. The overall phenomenon of symbol emergence within the group can be considered as a distributed effort to approximate $q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)$ without centralized control.
Mathematically, the CPC frames the process of symbol emergence in terms of social representation learning, analogous to the formation of internal representations by individual brains. As we introduced in Section 3, the representation learning and the inference of latent variables can be naturally extended to multi-agent reinforcement learning (RL) involving EmCom. Additionally, the formulation of CPC from the perspective of active inference can make us understand that this generative EmCom is not only for creating communication protocol but also for the environmental adaptation as a system of agents. Similarly, FEP and active inference are the general explanatory theory of the brain’s adaptation to the environment. In summary, the framework of generative EmCom based on the CPC hypothesis offers an explanation for language evolution to serve environmental adaptation at a collective level.
Generative EmCom offers an intriguing dual perspective, depending on whether one focuses on the collective or the individual. From a holistic perspective, the system appears to perform centralized representation learning of $w$ across all agents. Notably, $o^{k}$ can be easily extended to include multiple sensory modalities. From the perspective of the system overseer, or generative EmCom, multi-agent symbol emergence becomes analogous to multi-modal representation learning of external representations. Conversely, from the perspective of an individual agent, the system engages in multi-modal representation learning, influenced by and contributing to the inference of a shared latent variable $m$, which parameterizes the prior distribution common to all agents. The equivalence of these two interpretations is the cornerstone of the generative EmCom perspective. In essence, generative EmCom can be formulated as the representation learning of shared external representations, mediated through the representation learning of internal representations by each agent, based on multi-agent and multi-modal sensory-motor information.
TABLE I: Nomenclature and parameter details

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated a remarkable ability to capture extensive world knowledge, yet how this is achieved without direct sensorimotor experience remains a fundamental puzzle. This study proposes a novel theoretical solution by introducing the Collective World Model hypothesis. We argue that an LLM does not learn a world model from scratch; instead, it learns a statistical approximation of a collective world model that is already implicitly encoded in human language through a society-wide process of embodied, interactive sense-making. To formalize this process, we introduce generative emergent communication (Generative EmCom), a framework built on the Collective Predictive Coding (CPC). This framework models the emergence of language as a process of decentralized Bayesian inference over the internal states of multiple agents. We argue that this process effectively creates an encoder-decoder structure at a societal scale: human society collectively encodes its grounded, internal representations into language, and an LLM subsequently decodes these symbols to reconstruct a latent space that mirrors the structure of the original collective representations. This perspective provides a principled, mathematical explanation for how LLMs acquire their capabilities. The main contributions of this paper are: 1) the formalization of the Generative EmCom framework, clarifying its connection to world models and multi-agent reinforcement learning, and 2) its application to interpret LLMs, explaining phenomena such as distributional semantics as a natural consequence of representation reconstruction. This work provides a unified theory that bridges individual cognitive development, collective language evolution, and the foundations of large-scale AI.
keywords:
@@
emergent communication , large language model , world model , generative model , symbol emergence
[…]
1 Introduction
Large Language Models (LLMs) have achieved astonishing success, exhibiting a profound capacity for reasoning and knowledge retrieval across countless domains [12, 83]. A central puzzle, however, lies in their apparent grasp of the structure of the physical world. LLMs, trained on vast corpora of text, are not designed to interact with an environment and lack any direct sensorimotor experience [4, 10]. Yet, they seem to possess what could be described as a “world model.” How can an LLM acquire such a rich model of the world without ever having perceived or acted within it? This paper proposes a theoretical solution to this fundamental question.
The debate over whether LLMs truly possess world models is active and ongoing. A growing body of evidence suggests that LLMs acquire surprisingly rich representations of the real world. For instance, some studies show that LLMs implicitly learn representations corresponding to vision and hearing just by reading text [133], and that specific spatiotemporal representations emerge within their activations [36]. Furthermore, this internal knowledge proves functionally potent; LLMs can be used as world models for complex planning tasks [46], and their distributional semantics can be grounded in the physical world to guide robotic perception and action [93, 142]. Conversely, other work has revealed that LLMs are not reliable world simulators, suggesting their internal models are brittle or fundamentally different from those of embodied agents [136]. This conflicting evidence highlights a deep conceptual gap in our understanding.
To clarify our argument, we distinguish between two concepts of a world model [119, 26]:
  1. (1)
    Type 1 World Model (Internal Model): A subjective, internal model that an agent learns through its own sensorimotor interactions with the environment to predict future states and plan actions. This corresponds to the agent’s Umwelt [135].
  2. (2)
    Type 2 World Model (Model of the World): An objective, structured representation of knowledge about the world, its entities, and their relations, which may not be tied to a single agent’s direct experience.
We can frame this distinction as one between a subjective Type 1 World Model and an objective Type 2 World Model. Much of the confusion in current discourse arises from this distinction; while the term ”world model” in AI often refers to a Type 1 model, particularly in the context of model-based reinforcement learning and predictive coding [38, 40, 39, 35, 119], discussions about LLMs often imply they possess a Type 2 model. This presents a paradox: it is difficult to see how a disembodied LLM could form a subjective, embodied Type 1 model, yet it is equally unclear how it could acquire an objective Type 2 model without any access to the world.
This paper proposes a solution to this paradox. We argue that the seemingly objective knowledge within an LLM is, in fact, a pseudo-objective structure encoded in the distributional semantics of language. This structure emerges as a result of aggregating countless subjective, Type 1 world models from a society of embodied agents. From the perspective of generative emergent communication (Generative EmCom) an LLM does not model the world directly, but rather models the collective of these Type 1 models as externalized in language. This leads to our central hypothesis:
{mdframed}
[linewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt] The Collective World Model Hypothesis
Human language is not merely a communication protocol but serves as an externalized representation of a collective world model, which emerges from the decentralized, interactive sense-making processes of an entire society of embodied agents. LLMs acquire their world knowledge by learning a statistical approximation of this collective world model encoded in text corpora.
To provide a formal basis for this hypothesis, we introduce a new theoretical framework called Generative EmCom. This framework is built upon the Collective Predictive Coding (CPC) [117], which extends the principles of predictive coding, the free-energy principle (FEP) and the Bayesian brain hypothesis [50, 32, 34, 27] from individual brains to a societal level. Generative EmCom formalizes how a shared symbol system, i.e., language, emerges as multiple agents collectively seek to minimize their shared prediction errors about the world, a process we model as decentralized Bayesian inference.
However, such an integrative theory has been largely missing. The series of studies on emergent communication (EmCom) and symbol emergence has attempted to explain the formation of language [98, 68]. Yet, these approaches have often failed to bridge the gap between two interdependent aspects: first, the learning of an individual agent’s world model, which is grounded in its embodiment and environmental adaptation [35, 119]; and second, the collective emergence of a language that reflects this grounded knowledge. Our work aims to address this specific challenge by providing a unified framework.
This study, therefore, makes two primary contributions. First, we formalize the Generative EmCom framework, clarifying its relationship with conventional approaches and demonstrating its utility in multi-agent systems. Second, using this framework, we provide a principled, mathematical interpretation of LLMs as collective world models, offering a coherent explanation for their otherwise mysterious capabilities. This unified perspective bridges the gap between EmCom, cognitive development, and the foundations of large-scale AI, opening new frontiers for research. Indeed, the theoretical framework proposed herein has already begun to inspire new concrete models for multi-agent coordination in dynamic environments [88] and reward-independent communication in multi-agent reinforcement learning (MARL) [141].
The remainder of this paper is organized as follows: Section 2 reviews the current landscape of EmCom and world models, identifying the theoretical gap our work addresses. Section 3 details the theoretical framework of Generative EmCom. Section 4 explains how a collective world model can emerge from multi-agent interaction and cooperation. Section 5 applies this theory to reinterpret LLMs. Finally, Section 6 discusses the implications and limitations of our work, and Section 7 concludes the paper.
2 The Landscape of EmCom and World Models
@@
Language evolves and changes over time as a result of decentralized human communications [24, 98, 112, 113]. Sentences are generated to describe a wide range of phenomena, including external events, emotions, and intentions. In particular, the system of language is inherently dyanamics rather than static [137, 114, 126, 123]. As Peirce, the founder of semiotics, suggested, symbols, including language, can be characterized by a triadic relationship of sign, object, and interpretant [95, 96, 18]. Here, sign corresponds to words, sentences and other signals. In particular, the correspondence between sign and object, which is signified by a sign (i.e., signifier), is determined by an interpretant. In other words, the meaning of a sign, that is, language, depends on culture and context, and so on.
2.1 Conventional Approaches to EmCom and Their Limits
The study of how communication protocols and symbolic language emerge from multi-agent interaction, a field known as EmCom, has been explored through several major approaches, including language games, MARL, and iterated learning models (ILMs) [97, 68, 114, 146, 11]. A central theme in this research, particularly in studies employing language games and powered by deep neural networks [31, 84, 55], is the extent to which the resulting protocols, or emergent languages (EmLangs), resemble human language. For instance, questions have been raised about whether EmLangs exhibit compositionality [65, 14], follow well-known statistical properties of natural languages [15, 105, 131, 129], or can be shaped by cognitive constraints [104, 57].
A foundational paradigm for many of these investigations is the Lewis signaling game [73]. This approach, often aligned with Shannon’s information theory [110], can be regarded as a discriminative model for optimizing a communication protocol.
@@ 2.1 Conventional Approaches to EmCom and Their Limits
  3. 3.
We refer to $\mathcal{J}_{\textup{MI}}$ as the mutual information (MI)-maximizing objective function because it is known to be an evidence lower bound (ELBO) (up to constant) of the following mutual information between $X$ and $M$ [8, 99]:
This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively.
However, from the viewpoint of our paper, it is more fruitful to reinterpret these communication games through the lens of probabilistic generative models. This shift from a discriminative to a generative perspective reveals deeper connections to principles of cognition, learning, and the information-theoretic foundations of frameworks like the Variational Autoencoder (VAE) and the Information Bottleneck (IB) principle [3, 144].
This generative viewpoint reframes the signaling game as a problem of maximizing the ELBO [130]. The objective function becomes:
[…]
where $\beta$ is a hyperparameter. Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [45, 72, 66]. Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence. Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative frameworks discussed in later sections. All these formulations can be interpreted as representation learning with messages serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives. However, these signaling game-based approaches have inherent limitations: they typically assume a simple two-agent setting with asymmetric sender-receiver roles, and their extension to populated, decentralized settings is not straightforward.
Beyond these foundational models, the field of Semantic Communication (SC) has gained traction, aiming to transmit information based on meaning and effectiveness rather than bit-level fidelity [100]. This has spurred research into areas such as deep learning-based semantic encoders [138], contextual reasoning for shared understanding [109], and learning goal-oriented languages from interaction [29]. A comprehensive framework, Emergent Semantic Communication (ESC), seeks to integrate many of these threads by using causal reasoning in a neuro-symbolic architecture to create efficient emergent languages [125]. While these approaches advance the engineering of goal-oriented communication, they do not address how a society-wide, general-purpose language emerges to reflect a collective world model, which is the central question this paper addresses.
2.2 World Models for Individual Cognitive Agents
The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [37, 35, 119, 26]. The concept of world model has its origins in the early days of AI and robotics studies [86]. Initial studies on ML investigated techniques for agents to autonomously construct and adapt their world models [108, 116]. Currently, the term generally refers to predictive frameworks [39], predominantly implemented using deep neural network architectures.
The concept of world models is closely related to the idea of predictive coding (PC). PC is the idea that the brain constantly predicts sensory information and updates its internal models to minimize prediction errors. In the context of cognitive robotics and AI, world models provide the structure for representing and reasoning about the environment, whereas PC offers a mechanism for learning and updating these models based on sensory experiences. The free energy principle and active inference further unify these concepts, suggesting that both perception and action can be considered as processes of minimizing prediction errors or free energy. A theoretical connection exists between them [119] .
State transition: $\quad z_{t+1}\sim p(z_{t+1}|z_{t},a_{t})$ (5)
Observation: $\quad x_{t}\sim p(x_{t}|z_{t})$ (6)
Inference: $\quad z_{t}\sim q(z_{t}|x_{1:t},a_{1:t-1})$ (7)
This general framework has been instantiated in numerous influential models [26]. The “World Models” approach by Ha and Schmidhuber, for instance, demonstrated that an agent could learn a compressed latent dynamics model from pixel inputs and use it to train a controller entirely within its own “dream” [38, 37]. This line of research was significantly advanced by the Dreamer series of models, which successfully applied latent dynamics models to challenging continuous control tasks and even discrete Atari benchmarks and Minecraft [40, 39, 41, 42]. While many of these models rely on reconstructing observations, alternative approaches have also been explored. For example, some models focus on learning latent dynamics through contrastive learning without reconstruction [67, 89, 90], or by incorporating specific physical priors into the latent space, such as in NewtonianVAE [54, 91]. Furthermore, to handle information from multiple viewpoints, which is crucial in multi-agent settings, approaches like Multi-View Dreaming have been proposed to construct a world model from multi-view images using contrastive learning [63].
A prominent recent approach to learning such world models is the Joint-Embedding Predictive Architecture (JEPA), proposed by LeCun as a key component of a pathway towards autonomous machine intelligence [70]. Unlike generative approaches that attempt to predict missing information in pixel space, JEPA operates by predicting the representations of missing information in an abstract feature space [7]. This non-generative, self-supervised method aims to capture the underlying dependencies of the data, encouraging the model to learn semantic and predictive features rather than superficial details. The JEPA framework has proven effective for learning world models from various modalities, including images (I-JEPA) [7] and videos (V-JEPA) [9]. This line of research has recently culminated in V-JEPA 2, which extends the framework to enable robotic planning from video, further solidifying JEPA as a state-of-the-art method for an individual agent to develop its internal, predictive model of the world [6].
In summary, all the approaches discussed in this subsection, from classic POMDP-based models to modern architectures like JEPA, are fundamentally concerned with how a single agent learns a Type 1 World Model through its own interactions.
2.3 The Unaddressed Gap: From Individual Experience to Collective Language
The preceding review of research in EmCom (Section 2.1) and individual world models (Section 2.2) highlights a fundamental theoretical gap. On one hand, studies in EmCom have explored how agents can develop shared communication protocols, but these approaches often treat language as an abstract code, detached from the rich, embodied experiences that ground meaning in the real world. On the other hand, computational models and general theories that could provide a comprehensive and integrative understanding of symbol and language emergence [120, 123] are still lacking.
Specifically, these theories must address the critical interdependency between two aspects: first, the world modeling by individual agents, which is grounded in their embodiment and environmental adaptation [35, 119]; and second, the collective emergence of a language whose structure reflects this grounded world knowledge, for instance through distributional semantics [47, 36]. In essence, a crucial question remains unaddressed: What is the theoretical and computational mechanism that bridges the subjective, internal world of the individual with the objective, external world of collective language? Without a clear answer, our understanding of large-scale phenomena, such as the emergence of world knowledge in LLMs, remains incomplete. The need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [79, 123, 13, 117], and this paper aims to provide such a framework.
3 Generative EmCom: A Theoretical Framework
These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the FEP, PC, and world modeling. Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 35, 119]. This aligns with broader theoretical frameworks such as PC and the FEP. PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [50], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [94], which is associated with the idea of the Bayesian brain [27]. Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [64, 33, 20], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
@@ 3 Generative EmCom: A Theoretical Framework
Figure 1: (Left) probabilistic graphical model (PGM) representation of CPC in symbol emergence, that is, EmCom or language. The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system. The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system. (Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
3.1 The CPC Hypothesis
To address the challenges outlined in Section 2, this paper builds upon the CPC hypothesis [117, 121]. The idea extends the principles of PC and the FEP from the individual cognitive level to a societal level [122]. CPC posits that the emergence of symbol systems, particularly language, can be modeled as a decentralized Bayesian inference of a shared latent representation. It assumes that not only individual agents but also entire groups engaged in symbolic communication can be modeled as generative systems, aiming to minimize their collective free energy (CFE).
@@ 3.1 The CPC Hypothesis
A question is raised. How can we update the external representations, e.g., language, while our brains are disconnected physically? The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., Taniguchi et al. [124], Hagiwara et al. [44]). In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations. A representative example is Metropolis-Hastings Naming Game (MHNG) explained in Section 3.3. The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
@@
Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the CFE of a group of agents. This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
3.2 Formalizing Generative EmCom as Decentralized Bayesian Inference
The CPC hypothesis posits that a group of agents, interacting with each other and their environment, can be modeled as a single, large-scale probabilistic generative model. This allows us to formalize the emergence of a shared symbol system, such as language, as a process of decentralized Bayesian inference. For a group of $K$ agents, the generative and inference processes are defined as follows:
Generative Model: $\quad p(m,\{z_{k}\}_{k},\{x_{k}\}_{k}|\{a_{k}\}_{k})$
$=p(m)\prod_{k=1}^{K}p(x_{k}|z_{k},a_{k})p(z_{k}|m,a_{k})$ (8)
Inference Model: $\quad q(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})$
$=q(m|\{z_{k}\}_{k})\prod_{k=1}^{K}q(z_{k}|x_{k},a_{k})$ (9)
where for each agent $k$, $x_{k}$ is its observation, $a_{k}$ is its action, and $z_{k}$ is its internal representation. The variable $m$ represents the shared, external symbol system (i.e., message) that connects the agents. The inference process $q(z_{k}|x_{k},a_{k})$ corresponds to each agent’s individual representation learning, while $q(m|\{z_{k}\}_{k})$ corresponds to the collective process of symbol emergence, which can be instantiated by a language game as discussed in Section 3.3.
Under the free energy principle, the goal of this collective system is to minimize the variational free energy (VFE), which is equivalent to maximizing the ELBO of the log-likelihood of observations. We term this the CFE [122, 118]:
$F$ $=D_{\mathrm{KL}}\left[q(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})\parallel p(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})\right]$
$=\underbrace{D_{\mathrm{KL}}\left[q(m|\{z_{k}\}_{k})\|p(m)\right]}_{\text{Collective Regularization}}$
$\ \ \ +\sum_{k=1}^{K}\left(\underbrace{\mathbb{E}_{q}\left[-\log{p(x_{k}|z_{k},a_{k})}\right]}_{\text{Individual Prediction Error}}+\underbrace{D_{\mathrm{KL}}\left[q(z_{k}|x_{k},a_{k})\|p(z_{k}|m,a_{k})\right]}_{\text{Individual Regularization}}\right)$ (10)
This decomposition in Eq. (10) is central to our framework. It demonstrates that minimizing a single, unified objective function, the CFE, naturally gives rise to two distinct processes. The second term corresponds to each agent minimizing its own individual free energy, which involves learning to accurately reconstruct its observations (the prediction error term) while keeping its internal representations regularized (the individual regularization term). The first term, the Collective Regularization, drives the system to form a shared symbol $m$. This term quantifies the cost of encoding the collective internal states $\{z_{k}\}_{k}$ into the shared message $m$. The emergence of language is thus framed as a process that optimizes this trade-off between individual predictive accuracy and collective communicative efficiency.
This CFE formulation allows for a clear comparison with the ELBO-maximizing objective of signaling games presented in Section 2.1. While both are founded on the ELBO principle, they differ significantly in scope and structure. The signaling game ELBO typically models a two-agent, one-shot, and often asymmetric (sender/receiver) interaction. In contrast, the CFE provides a formulation for a population of $K$ agents engaged in a continuous, decentralized process of collective sense-making. The CFE’s explicit decomposition into individual and collective terms provides a richer, more structured model of the interplay between individual learning and social language emergence, a distinction that conventional signaling game formulations do not typically make explicit.
Table 1: Nomenclature and parameter details